# Computation of Robust Option Prices via Structured Multi-Marginal Martingale Optimal Transport

### v1

@@ 3.2 Formulation as a linear programming problem
We have now shown that the MOT problem (6) is equivalent to the OT problem (13), so that instead of computationally solving problem (6) we can solve problem (13), which has the sequential structure required for obtaining an efficient algorithm. In order to develop a computational method, we now consider the case when the support of the marginals is concentrated on a finite number of points — the objective function as well as the constraints of problem (13) can then be expressed in terms of tensors. We will now formalise this. With slight abuse of terminology, we refer to equation 13b as ‘martingale constraints’ from now on.
Suppose that the support of the given marginals $\{\mu_{t}\}_{t\in\mathcal{T}}$ of the price process is concentrated on a finite number of points. This is the situation for information given by a real-world financial market. Let the support of the given marginals be contained within the discrete set $\mathcal{S}\subset\mathbb{R}$ given by
$\mathcal{S}=\{s_{1},\dots,s_{n_{S}}\}$
where $n_{S}\in\mathbb{N}$ and $s_{1},\dots,s_{n_{S}}\in\mathbb{R}$. Let $s\in\mathbb{R}^{n_{S}}$ be the vector consisting of the elements in $\mathcal{S}$, that is
$s(i):=s_{i},\quad i=1,\dots,n_{S}.$
For each given marginal $\mu_{t}$ we define a vector $\upmu_{t}\in\mathbb{R}^{n_{S}}$ where each element is given by
$\upmu_{t}(i):=\mu_{t}\left(s(i)\right),\quad i=1,\dots,n_{S}.$ (16)
For practical convenience, we limit ourselves to only consider models $(\Omega,\mathcal{F},\mathbb{Q},S)$ where the support of $S_{t}$ is contained in $\mathcal{S}$ also for $t\in[T]\backslash\mathcal{T}$. Let $X$ be defined in terms of the price process $S$ as in equation 4 and let its (finite) range in $\mathbb{R}$ be denoted by $\mathcal{X}$. Then $\mathcal{X}$ is of the form
$\mathcal{X}=\{x_{1},\dots,x_{n_{X}}\}$
for some $n_{X}\in\mathbb{N}$ and $x_{1},\dots,x_{n_{X}}\in\mathbb{R}$. Let $x\in\mathbb{R}^{n_{X}}$ be the vector consisting of the elements of $\mathcal{X}$, that is
$x(i):=x_{i},\quad i=1,\dots,n_{X}.$
Note that depending on the choice of $h_{t}$, $t\in[T]$, we could have that $\mathcal{X}\equiv\mathcal{S}$ — this is the case for Examples 2.1 and 2.6.
Let us now take a model $(\Omega,\mathcal{F},\mathbb{Q},S)$ satisfying the above assumptions. For each $t\in[T]$ and almost all $\omega\in\Omega$ the price $S$ then takes one of the values represented by $s$,
$S_{t}(\omega)=s(\omega_{t}^{S}),$
for some $\omega_{t}^{S}\in\{1,\dots,n_{S}\}$. Note that $\omega_{t}^{S}$ denotes an index running from 1 to $n_{S}$. Also note that each time point $t$ is characterised by its own index so that in total we have $T+1$ indices $\omega_{0}^{S},\dots,\omega_{T}^{S}$, all running from 1 to $n_{S}$. Then the index tuple $(\omega_{0}^{S},\dots,\omega_{T}^{S})$ represents the outcome $\omega\in\Omega$. Similarly, the process $X$ at time $t\in[T]$ and for almost all $\omega\in\Omega$ takes one of the values represented by $x$,
$X_{t}(\omega)=x(\omega_{t}^{X})$
where $\omega_{t}^{X}\in\{1,\dots,n_{X}\}$. Here $\omega_{t}^{X}$ is an index running from 1 to $n_{X}$. By constructing a joint index
$\omega_{t}:=(\omega_{t}^{S},\omega_{t}^{X}),\quad t=0,\dots,T,$
a joint index tuple
$\boldsymbol{\omega}:=\left(\omega_{0},\dots,\omega_{T}\right)=\left((\omega_{0}^{S},\omega_{0}^{X}),\dots,(\omega_{T}^{S},\omega_{T}^{X})\right)$
and a joint index set
$\boldsymbol{\Omega}:=\big{(}\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}\big{)}^{T+1},$
we can identify each sample $\omega\in\Omega$ with an element $\boldsymbol{\omega}\in\boldsymbol{\Omega}$ via
$\bigg{(}\big{(}S_{0}\left(\omega\right),X_{0}\left(\omega\right)\big{)},\dots,\big{(}S_{T}\left(\omega\right),X_{T}\left(\omega\right)\big{)}\bigg{)}=\bigg{(}\big{(}s\left(\omega_{0}^{S}\right),x\left(\omega_{0}^{X}\right)\big{)},\dots,\big{(}s\left(\omega_{T}^{S}\right),x\left(\omega_{T}^{X}\right)\big{)}\bigg{)}.$
The index tuple $\boldsymbol{\omega}\in\boldsymbol{\Omega}$ thus represents one possible trajectory for the joint process $(S,X)$.
@@
We will now have a look at how the payoff can be represented within this framework. Let $\Phi:\boldsymbol{\Omega}\rightarrow\mathbb{R}$ be given by
$\Phi(\boldsymbol{\omega}):=\phi\left(s(\omega_{0}^{S}),\dots,s(\omega_{T}^{S})\right),\quad\boldsymbol{\omega}\in\boldsymbol{\Omega}.$
Then $\Phi(\boldsymbol{\omega})$ represents the payoff associated with the trajectory $\boldsymbol{\omega}\in\boldsymbol{\Omega}$. Also define $\Phi_{t}$ for $t\in[T]\backslash 0$ by
$\Phi_{t}(\omega_{t-1},\omega_{t}):=\phi_{t}\left(s(\omega_{t-1}^{S}),x(\omega_{t-1}^{X}),s(\omega_{t}^{S}),x(\omega_{t}^{X})\right),\quad\omega_{t-1},\omega_{t}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}.$
Then
$\Phi(\boldsymbol{\omega})=\sum_{t\in[T]\backslash 0}\Phi_{t}(\omega_{t-1},\omega_{t}),\quad\boldsymbol{\omega}\in\boldsymbol{\Omega},$ (17)
analogously to equation 5. In order to enforce a zero probability for every trajectory $\boldsymbol{\omega}\in\boldsymbol{\Omega}$ that does not respect equation 4, define a penalty term by
$I(\boldsymbol{\omega}):=\sum_{t\in[T]\backslash 0}I_{t}(\omega_{t-1},\omega_{t}),\quad\boldsymbol{\omega}\in\boldsymbol{\Omega},$
with
@@
$I_{t}(\omega_{t-1},\omega_{t}):=\begin{cases}0&\text{if }x(\omega_{t}^{X})=h_{t}\left(s(\omega_{t}^{S}),s(\omega_{t-1}^{S}),x(\omega_{t-1}^{X})\right)\\ \infty&\text{else}\end{cases}\\ \end{aligned},\quad t=2,\dots,T$
@@
$I_{1}(\omega_{0},\omega_{1}):=\begin{cases}0&\text{if }x(\omega_{0}^{X})=h_{0}\left(s(\omega_{0}^{S})\right)\text{ and }x(\omega_{1}^{X})=h_{1}\left(s(\omega_{1}^{S}),s(\omega_{0}^{S}),x(\omega_{0}^{X})\right)\\ \infty&\text{else}.\end{cases}$
By adding the penalty term $I$ to the payoff, we penalise forbidden trajectories $\boldsymbol{\omega}\in\boldsymbol{\Omega}$. Therefore, define $C_{t}$ for $t\in[T]\backslash 0$ by
$C_{t}(\omega_{t-1},\omega_{t}):=\Phi_{t}(\omega_{t-1},\omega_{t})+I_{t}(\omega_{t-1},\omega_{t}),\quad\omega_{t-1},\omega_{t}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\},$
and let $C(\boldsymbol{\omega}):=\sum_{t\in[T]\backslash 0}C_{t}(\omega_{t-1},\omega_{t})$ for $\boldsymbol{\omega}\in\boldsymbol{\Omega}$. Note that $C$ inherits the structure of $\Phi$ and $I$.
We have now presented the notation needed for formulating problem (13) as a linear programming problem; before doing so, we present some alternative notation that will simplify the expressions. By ordering the joint indices $\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$ according to
$(1,1),\dots,(n_{S},1),(1,2),\dots,(n_{S},2),{\dots}\hskip 2.84544pt{\dots},(1,n_{X}),\dots,(n_{S},n_{X}),$ (18)
$\Phi,I$ and $C$ can be identified with $T+1$-dimensional tensors in $\mathbb{R}^{(n_{S}n_{X})^{T+1}}$. Similarly $\Phi_{t},I_{t}$ and $C_{t}$, $t\in[T]\backslash 0$, can be represented by matrices in $\mathbb{R}^{n_{S}n_{X}\times n_{S}n_{X}}$. Define the matrix $\Delta\in\mathbb{R}^{n_{S}n_{X}\times n_{S}n_{X}}$ as
$\Delta(i,j):=(\mathbf{1}_{n_{X}}\otimes s)(j)-(\mathbf{1}_{n_{X}}\otimes s)(i),\quad i,j\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\},$ (19)
where it is assumed that the indices are ordered along each axis according to the rule given by (18). Recall the definitions of the projection operators $P_{t}$, for $t\in[T]$, and $P_{t_{1},t_{2}}$, for $t_{1},t_{2}\in[T]$. They can be identified with maps from $\mathbb{R}_{+}^{(n_{S}n_{X})^{T+1}}$ to $\mathbb{R}_{+}^{n_{S}n_{X}}$ and $\mathbb{R}_{+}^{n_{S}n_{X}\times n_{S}n_{X}}$, respectively, when the indices are ordered according to (18). Define $P^{S}:\mathbb{R}_{+}^{n_{S}n_{X}}\rightarrow\mathbb{R}_{+}^{n_{S}}$ as
$P^{S}(v)(\omega_{t}^{S}):=\sum_{\omega_{t}^{X}=1}^{n_{X}}v\left((\omega_{t}^{S},\omega_{t}^{X})\right),\quad\omega_{t}^{S}=1,\dots,n_{S}$
where $v\in\mathbb{R}_{+}^{n_{S}n_{X}}$ and let $P^{S}_{t}:=P^{S}\circ P_{t}$ for $t\in[T]$.
@@
$\underset{Q\in\mathbb{R}_{+}^{(n_{S}n_{X})^{T+1}}}{\min}\quad$ $\langle C,Q\rangle$ (20a)
subject to $\left(P_{t-1,t}(Q)\odot\Delta\right)\mathbf{1}_{n_{S}n_{X}}=\mathbf{0}_{n_{S}n_{X}},$ $t\in[T]\backslash 0$ (20b)
$P_{t}^{S}(Q)=\upmu_{t},$ $t\in\mathcal{T}.$ (20c)
We now show that problems (13) and (20) are equivalent.
@@
Suppose that we restrict problem (13) to models such that the support of the price process at each time point is contained within $\mathcal{S}$. Then problems (13) and (20) are equivalent.
Proof.
Let $(\Omega,\mathcal{F},\mathbb{Q},S)$ be a model feasible to problem (13) and note that we can define a probability measure $Q$ on $(\boldsymbol{\Omega},\sigma(\boldsymbol{\Omega}))$ by
$Q(\boldsymbol{\omega}):=\mathbb{Q}\left(\bigcap_{t=0}^{T}\bigg{\{}S_{t}=s(\omega_{t}^{S}),X_{t}=x(\omega_{t}^{X})\bigg{\}}\right),\quad\boldsymbol{\omega}\in\boldsymbol{\Omega}.$ (21)
It can be identified with a tensor, indexed by (18), in $\mathbb{R}_{+}^{(n_{S}n_{X})^{T+1}}$. Starting with the objective function, it holds that
$\mathbb{E}_{\mathbb{Q}}\left[\sum_{t\in[T]\backslash 0}\phi(S_{t-1},X_{t-1},S_{t},X_{t})\right]=\sum_{t\in[T]\backslash 0}\langle\Phi_{t},P_{t-1,t}(Q)\rangle=\langle\Phi,Q\rangle=\langle C,Q\rangle,$
where the last equality follows from the fact that $\langle I,Q\rangle=0$ for any tensor $Q$ defined as in equation 21. As for the marginal constraints (13c), it follows immediately from the definition of $Q$, for $t\in[T]$, that
$\mathbb{Q}\left(S_{t}=s(i)\right)=P_{t}^{S}(Q)(i),\quad i=1,\dots,n_{S}.$
Combining this elementwise with equation (16) and the constraint (13c) yields the constraint (20c). Similarly, for $t\in[T]\backslash 0$, it holds that
$\mathbb{E}_{\mathbb{Q}}\left[S_{t}|S_{t-1}=s(\omega_{t-1}^{S}),X_{t-1}=x(\omega_{t-1}^{X})\right]$ $=\frac{\big{(}P_{t-1,t}(Q)(\mathbf{1}_{n_{X}}\otimes s)\big{)}\left(\omega_{t-1}\right)}{P_{t-1}\left(Q\right)\left(\omega_{t-1}\right)}$
for $\omega_{t-1}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$, since $s(\omega_{t}^{S})=(\mathbf{1}_{n_{X}}\otimes s)(\omega_{t})$ when $\omega_{t}\in\{1,\dots,n_{S}\}\times\{1,\dots,n_{X}\}$ is ordered according to (18). By combining the martingale constraint (13b) elementwise with this we obtain
$P_{t-1,t}(Q)\left(\mathbf{1}_{n_{X}}\otimes s\right)=\left(\mathbf{1}_{n_{X}}\otimes s\right)\odot P_{t-1}\left(Q\right),\quad t=1,\dots,T.$
This is the martingale constraint (13b) formulated on vector form. By subtracting the right-hand side from the left-hand side and identifying the matrix $\Delta\in\mathbb{R}^{n_{S}n_{X}\times n_{S}n_{X}}$ the constraint (20b) follows.
Conversely, let $Q\in\mathbb{R}_{+}^{(n_{S}n_{X})^{T+1}}$ be feasible to problem (20). Then $(\boldsymbol{\Omega},\sigma(\boldsymbol{\Omega}),Q)$ is a probability space, on which we can identify $\Phi$ as a random variable. Assume that $Q$ is such that $\langle I,Q\rangle=0$. Then
$\langle C,Q\rangle=\langle\Phi,Q\rangle=\mathbb{E}_{Q}\left[\Phi\right],$
which means that the values of the objective functions (20a) and (13a) coincide. Also note that the model $(\boldsymbol{\Omega},\sigma(\boldsymbol{\Omega}),Q,s^{T+1})$, induced by the tensor $Q$ satisfying the constraints (20b) and (20c), satisfies the constraints (13b) and (13c). It remains to argue that there always exists a $Q$ such that $\langle I,Q\rangle=0$ whenever the feasible set of problem (20) is nonempty. Do so by noting that the feasible set of problem (13) is nonempty and that every feasible model $(\Omega,\mathcal{F},\mathbb{Q},S)$ induces $Q$ such that $\langle I,Q\rangle=0$. This completes the proof. ∎

### v2

@@ 3.2 Formulation as a linear programming problem
In order to develop a computational method, we now consider the case when the support of the given marginals is concentrated on a finite number of points — this is the situation for information given by a real-world financial market. For practical convenience, we limit ourselves to only consider models where the support of the intermediate marginals is also concentrated on a discrete set. The objective function as well as the constraints of problem (13) can then be expressed in terms of tensors; we will now formalise this. With a slight abuse of terminology, we refer to equation 13b as ‘martingale constraints’ from now on.
Let a family $\{\mathcal{S}_{t}\}_{t\in[T]}$ of discrete subsets of $\mathbb{R}$ be given by $\mathcal{S}_{t}=\{s_{t}^{1},\dots,s_{t}^{n^{\mathcal{S}}_{t}}\}$ with $n_{t}^{\mathcal{S}}\in\mathbb{N}$ and $t\in[T]$, where $\mathcal{S}_{t}$ for $t\in\mathcal{T}$ is the support of the given marginal $\mu_{t}$. We assume that the resulting grid $\mathcal{S}_{0}\times\dots\times\mathcal{S}_{T}$ is such that it is for $t\in[T]\backslash\mathcal{T}$ possible to construct intermediate marginals $\mu_{t}$, with support $\mathcal{S}_{t}$, respecting the convex order. Then the set of martingales, supported on $\mathcal{S}_{0}\times\dots\times\mathcal{S}_{T}$, that respects the given marginals is non-empty. Let for $t\in[T]$ the vector $s_{t}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ be the vector consisting of the elements in $\mathcal{S}_{t}$, that is $s_{t}(i):=s^{i}_{t}$ for $i\in\mathcal{I}^{\mathcal{S}}_{t}$, where $\mathcal{I}^{\mathcal{S}}_{t}:=\{1,\dots,n^{\mathcal{S}}_{t}\}$. For each given marginal $\mu_{t}$ we define a vector $m_{t}\in\mathbb{R}^{n_{t}^{\mathcal{S}}}$ where each element is given by
$m_{t}(i):=\mu_{t}\left(\{s_{t}(i)\}\right),\quad i\in\mathcal{I}_{t}^{\mathcal{S}},\quad t\in\mathcal{T}.$ (16)
Let for $t\in[T]\backslash 0$ the set $\mathcal{X}_{t}$ be such that $h_{t}(s_{t},s_{t-1},x_{t-1})\in\mathcal{X}_{t}$ for $(s_{t},s_{t-1},x_{t-1})\in\mathcal{S}_{t}\times\mathcal{S}_{t-1}\times\mathcal{X}_{t-1}$ and $\mathcal{X}_{0}$ be such that $h_{0}(s_{0})\in\mathcal{X}_{0}$ for $s_{0}\in\mathcal{S}_{0}$. Then each $\mathcal{X}_{t}$ is of the form $\mathcal{X}_{t}=\{x_{t}^{1},\dots,x_{t}^{n_{t}^{\mathcal{X}}}\}$ for some $n_{t}^{\mathcal{X}}\in\mathbb{N}$. Let for $t\in[T]$ the vector $x_{t}\in\mathbb{R}^{n_{t}^{\mathcal{X}}}$ be the vector consisting of the elements of $\mathcal{X}_{t}$, that is $x_{t}(i):=x_{t}^{i}$ for $i\in\mathcal{I}_{t}^{\mathcal{X}}$, where $\mathcal{I}_{t}^{\mathcal{X}}:=\{1,\dots,n_{t}^{\mathcal{X}}\}$. Note that depending on the choice of $h_{t}$, $t\in[T]$, we could have that $\mathcal{X}_{t}\equiv\mathcal{S}_{t}$ — this is the case for Examples 2.2 and 2.7. We emphasise that the joint state space $\mathcal{S}_{t}\times\mathcal{X}_{t}$ contains in total $n_{t}=n_{t}^{\mathcal{S}}n_{t}^{\mathcal{X}}$ states and is indexed by joint indices $(i_{t}^{\mathcal{S}},i_{t}^{\mathcal{X}})\in\mathcal{I}_{t}^{\mathcal{S}}\times\mathcal{I}_{t}^{\mathcal{X}}$. By letting $\mathcal{I}_{t}:=\{1,\dots,n_{t}\}$ and ordering the joint index set $\mathcal{I}_{t}^{\mathcal{S}}\times\mathcal{I}_{t}^{\mathcal{X}}$ according to
$(1,1),\dots,(n_{t}^{\mathcal{S}},1),(1,2),\dots,(n_{t}^{\mathcal{S}},2),{\dots}\hskip 2.84544pt{\dots},(1,n_{t}^{\mathcal{X}}),\dots,(n_{t}^{\mathcal{S}},n_{t}^{\mathcal{X}}),$ (17)
we can identify each element $i_{t}\in\mathcal{I}_{t}$ with an element of $(i_{t}^{\mathcal{S}},i_{t}^{\mathcal{X}})\in\mathcal{I}_{t}^{\mathcal{S}}\times\mathcal{I}_{t}^{\mathcal{X}}$; we will write $i_{t}$ for short for $i_{t}(i_{t}^{\mathcal{S}},i_{t}^{\mathcal{X}})$. Similarly, it is understood that $i_{t}^{\mathcal{S}}\in\mathcal{I}_{t}^{\mathcal{S}}$ and $i_{t}^{\mathcal{X}}\in\mathcal{I}_{t}^{\mathcal{X}}$ when $i_{t}(i_{t}^{\mathcal{S}},i_{t}^{\mathcal{X}})\in\mathcal{I}_{t}$. For notational convenience we also define $\mathcal{I}:={}_{t\in[T]}\mathcal{I}_{t}$, thus $(i_{0},\dots,i_{T})\in\mathcal{I}$ indicates that $i_{t}\in\mathcal{I}_{t}$ for $t\in[T]$.
@@
We will now see how the payoff can be represented on tensor form within this framework. Let $\Phi_{t}\in\mathbb{R}^{n_{t-1}\times n_{t}}$ for $t\in[T]\backslash 0$ be given by
$\Phi_{t}(i_{t-1},i_{t}):=\phi_{t}\big{(}s_{t-1}(i_{t-1}^{\mathcal{S}}),x_{t-1}(i_{t-1}^{\mathcal{X}}),s_{t}(i_{t}^{\mathcal{S}}),x_{t}(i_{t}^{\mathcal{X}})\big{)},\quad i_{t-1}\in\mathcal{I}_{t-1},i_{t}\in\mathcal{I}_{t}$
and define $\mathbf{\Phi}\in\mathbb{R}^{n_{0}\times\dots\times n_{T}}$ as $\mathbf{\Phi}(i_{0},\dots,i_{T}):=\sum_{t\in[T]\backslash 0}\Phi_{t}(i_{t-1},i_{t})$ for $(i_{0},\dots i_{T})\in\mathcal{I}$ analogously to equation 5. Then $\mathbf{\Phi}(i_{0},\dots,i_{T})=\phi(s_{0}(i_{0}^{\mathcal{S}}),\dots,s_{T}(i_{T}^{\mathcal{S}}))$ for $(i_{0},\dots,i_{T})\in\mathcal{I}$ and represents the payoff associated with the price evolution $(s_{0}(i_{0}^{\mathcal{S}}),x_{0}(i_{0}^{\mathcal{X}})),\dots,(s_{T}(i_{T}^{\mathcal{S}}),x_{T}(i_{T}^{\mathcal{X}}))$. In order to enforce a zero probability for every index tuple that does not respect equation 4, let $\mathbf{C}\in\mathbb{R}^{n_{0}\times\dots\times n_{T}}$ be a cost given by $\mathbf{C}(i_{0},\dots,i_{T}):=\sum_{t\in[T]\backslash 0}C_{t}(i_{t-1},i_{t})$ for $(i_{0},\dots,i_{T})\in\mathcal{I}$, with
@@
$C_{t}(i_{t-1},i_{t}):=\begin{cases}\Phi_{t}(i_{t-1},i_{t})&\text{if }x_{t}(i_{t}^{\mathcal{X}})=h_{t}\big{(}s_{t}(i_{t}^{\mathcal{S}}),s_{t-1}(i_{t-1}^{\mathcal{S}}),x_{t-1}(i_{t-1}^{\mathcal{X}})\big{)}\\ \infty&\text{else}\end{cases}\\ \end{aligned},\quad t=2,\dots,T$
@@
$C_{1}(i_{0},i_{1}):=\begin{cases}\Phi_{1}(i_{0},i_{1})&\text{if }x_{0}(i_{0}^{\mathcal{X}})=h_{0}\big{(}s_{0}(i_{0}^{\mathcal{S}})\big{)}\text{ and }x_{1}(i_{1}^{\mathcal{X}})=h_{1}\big{(}s_{1}(i_{1}^{\mathcal{S}}),s_{0}(i_{0}^{\mathcal{S}}),x_{0}(i_{0}^{\mathcal{X}})\big{)}\\ \infty&\text{else}\end{cases}$
being matrices in $\mathbb{R}^{n_{t-1}\times n_{t}}$ and $\mathbb{R}^{n_{0}\times n_{1}}$, respectively. By defining the cost $\mathbf{C}$ in this way, we penalise forbidden index tuples; note that $\langle\mathbf{C},\mathbf{Q}\rangle=\langle\mathbf{\Phi},\mathbf{Q}\rangle$ whenever $\langle\mathbf{C},\mathbf{Q}\rangle<\infty$. Also note that $\mathbf{C}$ inherits the structure of $\mathbf{\Phi}$ and therefore it decouples sequentially, as in equation 12.
We now present some notation needed for formulating problem (13) as a linear programming problem. Define for $t\in[T]\backslash 0$ the matrix $\Delta_{t}\in\mathbb{R}^{n_{t-1}\times n_{t}}$ as
$\Delta_{t}(i_{t-1},i_{t}):=s_{t}(i_{t}^{\mathcal{S}})-s_{t-1}(i_{t-1}^{\mathcal{S}}),\quad i_{t-1}\in\mathcal{I}_{t-1},i_{t}\in\mathcal{I}_{t},$ (18)
where we recall the connection between $i_{t}$ and $i_{t}^{\mathcal{S}}$ from (17). Recall the definitions of the projection operators $P_{t}$, for $t\in[T]$, and $P_{t_{1},t_{2}}$, for $t_{1},t_{2}\in[T]$ and note that they can be identified with maps from $\mathbb{R}_{+}^{n_{0}\times\dots\times n_{T}}$ to $\mathbb{R}_{+}^{n_{t}}$ and $\mathbb{R}_{+}^{n_{t_{1}}\times n_{t_{2}}}$, respectively. Then define for $t\in[T]$ a family of projection operators $P^{\mathcal{S}_{t}}:\mathbb{R}_{+}^{n_{t}}\rightarrow\mathbb{R}_{+}^{n_{t}^{\mathcal{S}}}$ via
$P^{\mathcal{S}_{t}}(m)(j):=\sum_{k\in\mathcal{I}_{t}^{\mathcal{X}}}m\big{(}i_{t}(j,k)\big{)},\quad j\in\mathcal{I}_{t}^{\mathcal{S}},\quad m\in\mathbb{R}_{+}^{n_{t}}$
— then $P^{\mathcal{S}_{t}}$ is the discrete analogue to integrating over the $\mathcal{X}_{t}$-component of a function (or distribution) on $\mathcal{S}_{t}\times\mathcal{X}_{t}$. Finally, let $P^{\mathcal{S}}_{t}:=P^{\mathcal{S}_{t}}\circ P_{t}$ for $t\in[T]$.
@@
$\underset{\mathbf{Q}\in\mathbb{R}_{+}^{n_{0}\times\dots\times n_{T}}}{\min}\quad$ $\langle\mathbf{C},\mathbf{Q}\rangle$ (19a)
subject to $\left(P_{t-1,t}(\mathbf{Q})\odot\Delta_{t}\right)1_{n_{t}}=0_{n_{t-1}},$ $t\in[T]\backslash 0$ (19b)
$P_{t}^{\mathcal{S}}(\mathbf{Q})=m_{t},$ $t\in\mathcal{T}.$ (19c)
Note that it has an optimal solution since its feasible set is bounded and non-empty, but that it may not be unique. The next result shows that that problems (13) and (19) are equivalent — the proof is deferred to Appendix: Additional proofs.
@@
Suppose that we restrict problem (13) to models such that the support of the price process at each time point $t\in[T]$ is $\mathcal{S}_{t}$. Then problems (13) and (19) are equivalent.